### v1

# Affine Frequency Division Multiplexing (AFDM) for 6G: Properties, Features, and Challenges

Abstract
Affine frequency division multiplexing (AFDM) is an emerging waveform candidate for future sixth generation (6G) systems offering a range of promising features, such as enhanced robustness in heterogeneous and high-mobility environments, as well as inherent suitability for integrated sensing and communications (ISAC) applications. In addition, unlike other candidates such as orthogonal time-frequency space (OTFS) modulation, AFDM provides several unique advantages that strengthen its relevance to practical deployment and standardization in 6G. Notably, as a natural generalization of orthogonal frequency division multiplexing (OFDM), strong backward compatibility with existing conventional systems is guaranteed, while also offering novel possibilities in waveform design, for example to enable physical-layer security through its inherent chirp parametrization. In all, this article provides an overview of AFDM, emphasizing its suitability as a candidate waveform for 6G standardization. First, we provide a concise introduction to the fundamental properties and unique characteristics of AFDM, followed by highlights of its advantageous features, and finally a discussion of its potential and challenges in 6G standardization efforts and representative requirements.
Index Terms:
Affine frequency division multiplexing (AFDM), sixth generation (6G), standardization, waveform design, delay-Doppler channel, integrated sensing and communications (ISAC).
I Introduction
The telecommunications community, including academia and industry, is working vigorously in the development of sixth generation (6G) wireless systems expected to be standardized by 2030 [2]. The objective for 6G is to deliver transformative capabilities, including peak data rates up to 1 Terabits-per-second (Tbps), ultra-low latency below 1 millisecond (ms), and massive connectivity supporting up to $10^{7}$ devices per km2. These ambitious goals are driven by the six well-known IMT-2030 usage scenarios [3], including integrated sensing and communications (ISAC), ubiquitous connectivity, and immersive communication, which are complemented by selected use cases such as Internet-of-Things (IoT), vehicle-to-everything (V2X), space-air-ground integrated networks (SAGINs), and digital twin networks (DTNs) [4].
Enabling such advanced applications and satisfying such stringent requirements demand, however, wireless systems to operate in increasingly higher frequency bands, such as the millimeter wave (mmWave) and Terahertz (THz) spectra, subject to multipath and high-mobility. This gives rise to doubly-dispersive channels that exhibit both time- and frequency-selectivity [5], leading to severe deterioration of received signals and, consequently, substantial challenges for communications systems employing conventional waveforms [6]. In particular, the orthogonal frequency division multiplexing (OFDM) waveform, widely used in fourth generation (4G) and fifth generation (5G) systems, struggles in such environments due to severe inter-symbol interference (ISI) and inter-carrier interference (ICI) caused by multipath delays and especially Doppler shifts, which disrupts OFDM subcarrier orthogonality.
Furthermore, 6G is anticipated to bring a paradigm shift towards a native, and hence, optimized, support of ISAC [2], enabling a wide range of multifunctional services, real-time situational awareness, improved spectrum utilization, and high-resolution localization [7]. These features are not easily available through conventional waveforms, which are susceptible to Doppler interference and prevent accurate velocity estimation,estimation, thus motivating the design of novel waveforms that can offer strong Doppler robustness and high-resolution sensing capabilities.
Given the above, innovative waveform design for 6G has gained great interest in recent years. In particular, in order to address the aforementioned challenges, researchers have explored various alternatives to maintain subcarrier orthogonality in doubly-dispersive conditions. Examples of waveforms that meet such requirements are: orthogonal time-frequency space (OTFS), whereby data is modulated in the delay-Doppler (as opposed to time-frequency) domain, improving robustness against Doppler fading [8]; and orthogonal chirp division multiplexing (OCDM), which leverages chirp-shaped subcarriers for enhanced performance in dispersive channels [9].
Another recent proposal is affine frequency division multiplexing (AFDM), which has emerged as a promising candidate for 6G [10], whereby data based is modulated via the affine Fourier transform (AFT), an affine generalization of the conventional linear Fourier transform (FT) employed in OFDM. By leveraging two tunable parameters that control the twistedness of the affine time-frequency domain, AFDM is capable of adapting to the specific characteristics of doubly-dispersive channels, enabling optimal diversity gain and delay-Doppler orthogonality. In addition, these tunable parameters facilitate a wide range of waveform design opportunities, supporting inherent advanced functionalities, such as index modulation (IM), peak-to-average power ratio (PAPR) reduction, and physical-layer security [11].
Within ongoing discussions on candidate waveforms for 6G, a key point of debate among industry, academia, and standardization bodies such as the 3rd Generation Partnership Project (3GPP), is whether 6G necessitates entirely new waveform paradigms or refinements of existing ones. Rather than replacing legacy waveforms like OFDM, many advocate for expanding the waveform toolbox to enable flexible, context-aware operation, highlighting the need for candidates that balance performance, compatibility, and deployability. In this context, AFDM stands out for its inherent compatibility with OFDM-based systems, enabled by its generalization of the FT to the AFT, facilitating seamless integration and coexistence with existing 4G and 5G infrastructure [12].
All in all, the above collectively set AFDM as a strong contender for standardization in the 6G era. Therefore, we aim at providing a comprehensive tutorial on, and motivation for AFDM, exploring its fundamental properties and distinctive features, as well as challenges and opportunities for its adoption in 6G networks. The article is structured as follows: Section II details the fundamentals and key properties of the AFDM waveform; Section III examines eight of the advantageous features of AFDM and evaluates their implications to relevant 6G applications; Section IV then addresses another set of eight points on opportunities, challenges, and pathways towards standardization of AFDM in future 6G systems; and finally, Section V concludes with a summary and prospects.
II Fundamentals and Key Properties
Figure 1: Illustrative comparison of OFDM, OCDM, AFDM, and OTFS waveforms in terms of: a) the modulator structure in terms of the core Fourier transform and the relationship of the chirp frequencies $c_{1}$ and $c_{2}$, which shows that AFDM generalizes OFDM and OCDM, b) the time-domain subcarriers (real component only), c) the time-frequency occupancy of the subcarriers, d) effective channel structure with four arbitrary integer delay-Doppler taps, and e) the same channels of d) with fractional Doppler shifts, causing Doppler-domain leakage.
In this section, we briefly introduce the fundamentals of the AFDM signal modulation, highlighting the inherent properties of the waveform, providing a background towards the consequent features and applications to be elaborated in the following sections. Given the breadth of the audience and the tutorial nature of the IEEE Magazine, this article presents an illustrative and intuitive overview of the AFDM fundamentals, omitting detailed derivations. Readers interested in the full theoretical formulation and rigorous analysis of the AFDM are encouraged to consult [10, 6].
II-A Signal Generation and Detection
At each transmission instance, a block of $N$ information symbols is mapped onto $N$ subcarriers through the inverse discrete affine Fourier transform (IDAFT), which serves as the modulation kernel of AFDM. Unlike the conventional notion of static, sinusoidal subcarriers in OFDM (i.e., pure-tone), which maintain orthogonality in the frequency domain, AFDM employs chirp-based subcarriers that sweep across the frequency bandwidth over time, as illustrated in the subcarrier visualizations of Fig. 1. Note that despite their time-varying nature, the $N$ chirp subcarriers maintain mutual orthogonality across both time and frequency domains, forming a robust basis well-suited for time-varying propagation environments.
Furthermore, the frequency sweep rate of the chirp subcarriers in AFDM is fully configurable via the two tunable chirp parameters inherent to the IDAFT (often denoted by $c_{1}$ and $c_{2}$). As will be further elaborated upon later, this unique design flexibility through the inclusion of tunable chirp parameters - not present in contending waveforms - is a key feature of AFDM, which ensures orthogonality and full diversity in doubly-dispersive channels, in addition to enabling a plethora of waveform optimization and applications.
Trivially, the IDAFT of AFDM parallels the role of the inverse discrete Fourier transform (IDFT) modulator in OFDM, and also stands in contrast to other chirp-subcarrier based waveforms such as OCDM leveraging the inverse discrete Fresnel transform (IDFnT) for modulation, in which the freqeuncy sweep rate is fixed and not configurable.
Considering the implementation of the modulator, the core IDAFT can be efficiently be decomposed into a chirp-domain pre-processing operation, followed by a standard IDFT, and a chirp-domain post-processing stage, as visualized in Fig. 1. This structure enables efficient implementation using the conventional fast Fourier transform (FFT) technology core to the 4G and 5G architectures. Furthermore, recent works have even demonstrated that both chirp pre-/post-processing steps be implemented using only discrete Fourier transform (DFT)/IDFT operations, consolidating the overall computational efficiency and compatibility of the AFDM modulator.
Finally, before transmission, the AFDM signal must be prepended with a chirp-periodic prefix (CPP), which while structurally similar to the conventional cyclic prefix (CP) in OFDM, includes deterministic phase adjustments determined by the chirp parameters. With the CPP, AFDM satisfies the quasi-periodicity conditions and ensures the circular convolution compatibility to preserve the symbol orthogonality.
At the receiver, demodulation is performed using a matched forward discrete affine Fourier transform (DAFT), which is configured with the same chirp parameters $c_{1}$ and $c_{2}$ used in the transmitter-side IDAFT, ensuring coherent symbol recovery. For appropriately selected chirp parameters, whose optimal values can be obtained in closed-form based on the channel statistics (maximum unambiguous delay and Doppler shift), and a desired guard width, the effective channel matrix becomes a structured superposition of shifted band-diagonal components. Each such component corresponds to a distinct delay-Doppler path, and their orthogonality minimises ISI and ICI even under significant channel dispersion, further ensured by the selected guard width parameter.
II-B Affine Fourier Domain and the Evolution of 6G Waveforms
Chirp signals are well-known to have favorable properties under both time shifts (delays) and frequency shifts (Doppler), which intuitively suggests them to be well-matched to channels where such effects are jointly present, i.e., the doubly-dispersive channels. As described in the previous section, AFDM exploits this notion, and modulates the data symbols unto the orthogonal chirp subcarriers via the inverse discrete affine Fourier transform (IDAFT).
Trivially, the IDAFT originates from the affine Fourier transform (AFT), also referred to as the linear canonical transform (LCT) [13], a family of unitary linear operators defined by four configurable parameters $(a,b,c,d)$, satisfying $ad-bc=1$. By selecting appropriate values of the parameters $a,b,c,$ and $d$, the AFT generalizes a wide range of well-known and fundamental transforms as its special cases.
For example, the AFT is reduced to the conventional Fourier transform (FT)with fixed parameters $(a,b,c,d)=(0,1,-1,0)$, while the fractional Fourier transform (FrFT)generalizes this by introducing a tunable angle parameter $\alpha\in[0,2\pi)$ such that $(a,b,c,d)=(\cos\alpha,\sin\alpha,-\sin\alpha,\cos\alpha)$. On the other hand, the Fresnel transform (FnT), associated with free-space wave propagation and used in the modulation of OCDM, is obtained when $(a,b,c,d)=(1,z,0,1)$, where $z$ is a tunable parameter representing the propagation distance.
Furthermore, complex transforms including the Laplace transform (LT) can also be obtained, by setting the parameters as $(a,b,c,d)=(0,\sqrt{-1},\sqrt{-1},0)=(0,j,j,0)$, but such complex transforms are non-unitary and lack an orthonormal basis, rendering them unsuitable for energy-preserving signal modulation and orthogonal multiplexing for waveform design.
[…]
III AFDM: Features and Applications
III-A Delay-Doppler Robustness and Diversity
One of the most critical requirements for 6G waveforms is the robustness to delay and Doppler dispersion - i.e., the ability to preserve orthogonality and full diversity in presence of rich multipath and Doppler shifts. This becomes more prominent in the envisioned high-frequency and dynamic environments, including urban V2X and SAGIN architectures.
The AFDM waveform stands out in this regard over other waveforms in the communications performance, as illlustrated in Fig. 2, offering inherent robustness to doubly-dispersive channels with arbitrary delay-Doppler statistics. Through the parametrizable chirp subcarriers, AFDM has been shown to guarantee full diversity without the need for precoding or equalization enhancements, by optimizing the first chirp parameter $c_{1}$, for which a closed-form solution is available [10].
This contrasts with conventional OFDM, where each symbol is confined to a single time-frequency subcarrier, resulting in a maximum diversity order of one in doubly-dispersive channels. While OTFS improves over the OFDM by spreading each symbol over the entire time-frequency grid via its delay-Doppler representation and transforms, uncoded OTFS has been shown to fall short of full diversity in the high-signal-to-noise ratio (SNR) regime unless mitigation strategies such as delay-Doppler precoding are applied [14].
Moreover, effective channel analysis has revealed that AFDM is inherently more robust fractional Doppler interference compared to other waveforms [10, 6], leveraging a tunable guard parameter to provide additional protection between each delay-Doppler tap.
Figure 2: BER performance of OFDM, OCDM, OTFS, and AFDM over a doubly-dispersive channel with $P=3$ paths. System parameters are: number of subcarriers $N=64$, constellation cardinality $M=4$, OTFS grid size $K\times L=8\times 8$, carrier frequency $f_{c}=50$GHz, bandwidth $B=150$MHz, AFDM guard parameter $\xi=1$, AFDM pre-chirp parameter $c_{2}=1/{2\pi N}$.
III-B Pilot Design and Channel Estimation
Accurate and efficient channel estimation is fundamentally crucial for reliable wireless communications, where the consideration of pilot placement also is necessary in balancing estimation accuracy, spectral efficiency, and system latency.
In terms of pilotting efficiency, AFDM benefits from the inherent sparsity and full delay-Doppler representation enabled by its DAFT-domain formulation. A single well-placed pilot symbol, separated from data by only a small one-dimensional guarding determined by the maximum channel delay, can capture the entire delay-Doppler profile without data interference.
This stands in contrast to OFDM, which requires multiple pilot tones scattered across time and frequency, and OTFS, which requires larger two-dimensional guard symbols around the pilot to ensure data interference-free estimation. By eliminating the need for such extensive pilot resources, AFDM allows for tighter pilot-data packing, enhancing spectral efficiency and reducing overhead, which is particularly beneficial in latency- and bandwidth-constrained scenarios.
Building on this efficient pilot design, it has been shown that the full delay-Doppler channel in AFDM can be efficiently estimated using a single pilot symbol and low-complexity methods [10], thanks to the sparse effective channel structure composed of multiple band-diagonal components, unlike the block-shifted diagonal structure of OTFS or the interference-prone diagonal structure of OFDM [6], as illustrated in Fig. 1.
III-C CSI-based Radar Parameter Estimation
Fundamentally, the full delay-Doppler representation of the doubly-dispersive channel is very closely linked to the radar parameters of environmental scatterers. Namely, each multipath component is associated with its delay and Doppler shift, which correspond to the bistatic range and velocity of the scatterer (or target). This duality has motivated a wide range of methods for extracting such physical parameters from the estimated channel state information (CSI), or conversely to perform channel estimation directly via parameterized models that exploit the deterministic structure of the delay-Doppler channel, which has shown great success and efficiency [15].
Such estimation techniques can be categorized into two main approaches: on-grid and off-grid methods. On-grid estimation involves discretizing the delay-Doppler space based on assumed maximum delay and Doppler values, followed by a search over this space, using either maximum likelihood estimation or compressive sensing-based sparse recovery. In contrast, off-grid approaches aim to circumvent discretization and perform estimation directly on the continuous delay-Doppler parameters, leveraging various techniques such as basis expansion methods and iterative algorithms.
III-D Conventional Radar Processing and Ambiguity Function
While CSI-based radar parameter extraction has become central to many modern communication-centric sensing frameworks, it is equally critical to consider the classical signal processing foundation of radar sensing: correlation-based matched filtering.
This classical perspective is especially relevant for AFDM, whose underlying chirp subcarriers strongly resemble the gold standard of chirp structures historically used in radar systems. In this context, the ambiguity function (AF) is fundamental to quantify how effectively a probing waveform can resolve targets in the delay and Doppler dimensions, characterizing the resolution and sidelobe suppression.
As illustrated in Fig. 3, in contrast to the poor ambiguity properties of OFDM in the Doppler domain, suffering from severe Doppler smearing and limited velocity resolution, or similarly of the OCDM in the delay-domain, AFDM exhibits a sharply localized, spike-like ambiguity function in both of the delay and Doppler dimensions, similarly to the OTFS, but has additionally shown to provide delay-Doppler lattice alignment options via the configurable of the chirp parameters.
(a) Zero-Doppler cut.
(b) Zero-delay cut.
Figure 3: Normalized ambiguity functions of OFDM, OCDM, OTFS, and AFDM in the delay-Doppler domain, with unitary sensing symbols. System parameters are: number of subcarriers $N=64$, OTFS grid size $K\times L=8\times 8$, carrier frequency $f_{c}=50$GHz, bandwidth $B=150$MHz, sampling frequency $f_{s}=2B$ Hz, AFDM guard parameter $\xi=1$, AFDM pre-chirp parameter $c_{2}=1/{2\pi N}$.
III-E AFDM for Integrated Sensing and Communciations (ISAC)
The above features of AFDM, particularly the excellent communications and radar ambiguity properties in doubly-dispersive channels, make it well-suited also for communication-centric ISAC, where the communication waveform itself is leveraged to extract high-resolution environmental information without requiring dedicated sensing signals, while also jointly supporting high-rate data transmission.
Indeed, recent works have shown that AFDM achieves accurate delay and Doppler parameter recovery with minimal sensing overhead but without sacrificing communication performance, leveraging embedded pilot modulation, subcarrier-level index modulation, and bilinear estimation methods for joint channel and data recovery [15].
III-F Frequency-Domain Processing
As expected from the close relationship between the AFDM and OFDM, for specific values of the chirp parameters $c_{1}$ and $c_{2}$, and certain subcarrier lengths $N$ still achieving full diversity and subcarrier orthogonality - closed-form frequency-domain expressions of the AFDM waveform are available, enabling straightforward implementation and conventional analysis [12]. In other words, AFDM can be viewed as a precoded OFDM system, where the precoder maps the data from the DAFT domain into the frequency domain.
This compatibility with the frequency-domain processing of OFDM is a key advantage of AFDM, as it allows for the direct application of established OFDM techniques, such as frequency-domain pilot placement, channel estimation and equalization, and digital beamforming.
Furthermore, this also carries an important implication that a smooth transition from 5G to 6G, i.e., from OFDM to AFDM, can be achieved without major modifications to existing hardware and infrastructure. This makes AFDM a practical and backward-compatible candidate for future wireless systems, as further discussed in Section IV-A.
III-G Modulation Flexibility and Novel Applications
The tunable parameters of AFDM, $c_{1}$ and $c_{2}$, offer design flexibility absent in conventional waveforms like OFDM and OTFS. While $c_{1}$ must satisfy a minimal condition to ensure full diversity, $c_{2}$ remains largely unconstrained and can be tuned freely. Importantly, such configurations preserve the core benefits of AFDM, including diversity, pilot design simplicity, and favorable ambiguity characteristics.
This flexibility enables secondary functionalities such as index modulation via discrete $c_{2}$ sets or chirp permutations, and also enables physical-layer security features (see following section). These make AFDM especially suitable for systems requiring both high performance and embedded functionality within a unified, standard-compatible waveform.
III-H Inherent Physical-Layer Security
AFDM provides inherent physical-layer security through configurable the chirp parameters ($c_{1}$, $c_{2}$) and transform-domain methods such as chirp permutation [11], without degrading performance.
Legitimate users can treat these configurations as shared secret keys, while unauthorized receivers face infeasible brute-force search over a continuous parameter space, shown to be infeasible even under quantum-accelerated attackes. By contrast, OFDM employs fixed DFT-based subcarriers with no intrinsic secrecy, and while OTFS offers limited structural confidentiality via discrete grid size options, but the finite configuration space is constrained by channel diversity requirements. AFDM, with its broader and continuous parameter space, enables significantly stronger waveform-level confidentiality.
IV Standardization Considerations for 6G:
Challenges, Opportunities, and Pathways
TABLE I: Mapping of AFDM properties and features (Sections II-III) to relevant 6G standardization considerations (Section IV).
Fundamental Properties and Features of AFDM Waveform (Sections II-III)
Orthogonal Chirp-Domain
Subcarriers
(II-A)
Generality of
Affine Fourier Transform
(II-B)
Delay-Doppler Robustness
& Full Diversity
(III-A)
Pilot Design
& Channel Estimation
(III-B)
CSI-based Radar
Parameter Estimation
(III-C)
Radar Processing
& Ambiguity Function
(III-D)
AFDM for Integrated Sensing and Commun.
(III-E)
Frequency-Domain
Processing
(III-F)
Modulation Flexibility
& Novel Applications
(III-G)
Inherent
Physical-Layer Security
(III-H)
Standardization Considerations for 6G
(Section IV)
A. Backward Compatibility with the OFDM Standard
B. Multi-RAT Coexistence and Spectrum Sharing
C. PAPR and Hardware Efficiency
D. Ultra-Low Latency: Implementation Complexity
E. Standardized Parameter Configuration
F. Integration with Emerging Use Cases
G. Robustness to Practical Impairments
H. Testbeds, Trials, and Reference Designs
The preceding sections have provided a comprehensive overview of the AFDM waveform, outlining its key signal-theoretic features and practical advantages across multiple domains, including channel resilience, diversity, spectral properties, and computational efficiency. Building upon this foundation, this section explains how these technical features align with some emerging and current 6G requirements and design principles outlined by standardization bodies and academia, such as 3GPP [2], International Telecommunication Union - Radiocommunication Sector (ITU-R) [3], and Institute of Electrical and Electronics Engineers (IEEE).
In this context, this section highlights the challenges and opportunities associated with integrating AFDM into the evolving 6G ecosystem, offering a roadmap that connects physical layer innovations and features of the waveform, with the practical deployment pathways, as summarized in Table I, and also qualitatively compared in terms of the various candidate waveforms in Fig. 4.
Figure 4: Qualitative comparison of the features and standardization-related properties of various waveforms: OFDM, OCDM, OTFS, and AFDM, which illustrates that AFDM is the best all-rounder waveform current under consideration for ISAC-enabling 6G systems, without significant sacrifice compared to OTFS and OFDM in terms of their specific advantages in hardware efficiency and low latency.
IV-A Backward Compatibility with the OFDM Standard
As depicted in Fig. 1, AFDM can be interpreted as a generalization of both OFDM and OCDM. Furthermore, as identified in Section III-F since frequency-domain processing techniques developed for OFDM are directly applicable to AFDM, the transition to AFDM can be achieved with minimal cost by leveraging existing OFDM infrastructure.
In other words, upgrading an OFDM system to AFDM merely involves the inclusion of a precoding block at the transmitter and a corresponding decoding block at the receiver. This modification is particularly cost-effective in software-defined radio platforms, and enables the seamless reuse of a wide range of components and protocols already specified in standards such as those from 3GPP and IEEE. As a result, AFDM supports a cost-efficient migration path from 5G to 6G, facilitating incremental upgrades and hybrid deployments during early stages of 6G adoption.
While similar compatibility arguments can also be made for OTFS, AFDM exhibits greater flexibility and robustness, enhancing its backward compatibility with OFDM. As previously discussed, AFDM achieves full diversity for any number of subcarriers $N$, depending solely on the chirp parameters $c_{1}$ and $c_{2}$. In contrast, full diversity in OTFS is only possible for certain divisors of $N$, a constraint that becomes particularly limiting in narrowband applications.
This distinction is especially relevant in use cases such as narrowband (NB)-IoT, where the number of subcarriers may be as small as $N=12$, corresponding to a single resource block defined in the 3GPP specification. In such scenarios, only AFDM can simultaneously ensure backward compatibility with OFDM and robustness to doubly dispersive channels, whereas OTFS falls short in this regard.
IV-B Multi-RAT Coexistence and Spectrum Sharing
As the wireless landscape becomes increasingly heterogeneous, waveform compatibility across multiple radio access technology (RAT) (multi-RAT) is essential for enabling seamless coexistence and efficient spectrum utilization. The previously elaborated compatibility of AFDM with OFDM, provides a strong foundation for cost-effective migration from 5G to 6G, specifically in facilitating hybrid deployments where multiple waveforms share spectrum or infrastructure.
In particular, the spectral and structural similarity between AFDM and OFDM allows for flexible spectrum partitioning and orthogonal resource allocation, enabling both waveforms to operate side-by-side with minimal interference - as shown by AFDM reducing to the OFDM with appropriate chirp parameters. This is especially relevant for legacy support, where incumbent 5G services must coexist with new 6G features during the early stages of deployment. These attributes support the vision of future networks where spectrum resources are flexibly and efficiently allocated across diverse services and technologies across generations, in alignment with evolving ITU-R and 3GPP spectrum sharing frameworks.
IV-C PAPR and Hardware Efficiency
Hardware efficiency is a key consideration for waveform adoption in future wireless standards. PAPR, in particular, remains a critical issue, as high values reduce power amplifier efficiency and raise hardware costs, especially in mobile and battery-powered 6G devices. Therefore, while AFDM offers strong delay-Doppler performance, maintaining favorable peak-to-average power ratio (PAPR)characteristics is still essential for practical deployment.
To address this issue, techniques such as chirp parameter tuning and DAFT-domain spreading have been proposed specifically for AFDM, providing design flexibility without degrading performance in doubly-dispersive channels. Additionally, the compatibility with OFDM systems also enables the reuse of established PAPR reduction methods, including tone reservation, windowing, and pulse shaping.
Furthermore, as will be elaborated in the following, compared to waveforms like OTFS, AFDM has been shown to achieve a lower-complexity modulator structure, reducing hardware overhead and improving energy efficiency across diverse 6G use cases, from high-speed access to massive machine-type communications.
IV-D Ultra-Low Latency: Implementation Complexity
The practical adoption of any new waveform hinges on its ability to support real-time processing under stringent latency and complexity constraints, especially in 6G systems that demand extremely low latency and high throughput. Although AFDM relies on the advanced DAFT, it was shown - as discussed in previous sections - to be implemented efficiently using structured fast transform algorithms, remaining tractable for even larger systems increasingly expected in 6G networks.
In addition to low-complexity modulation and demodulation, AFDM also benefits significantly from the sparse and structured effective channel response, enabling simplified equalization and efficient pilot design. Compared to OTFS, which requires a two-dimensional piloting guard for effective channel estimation and typically more complex signal processing for channel reconstruction, AFDM offers a more streamlined receiver architecture. These advantages position AFDM as a waveform that can increasing meet real-time requirements across a wide range of devices, from high-performance base stations to resource-constrained user terminals.
IV-E Standardized Parameter Configuration
As discussed, one of the defining feature of AFDM is its flexible parametrization of continuous-valued chirp frequencies ($c_{1}$, $c_{2}$), enabling adaptation to channel conditions, system requirements, and additional applications. However, this flexibility will also necessitate standardized procedures for parameter sharing and signaling. Unlike OFDM, where subcarrier spacing and symbol duration are fixed by the numerology, AFDM supports a broader configuration space. Therefore, to ensure interoperability and consistent performance across vendors and deployments, standards must define permissible values or ranges for the chirp parameters, potentially based on link type, mobility profile, or application class.
Novel signaling mechanisms must also be considered to convey the chirp configurations in both uplink and downlink, covering various components such as initial access, adaptive reconfiguration during handover, and multi-cell coordination. Well-defined parameter profiles and negotiation protocols will be essential to minimize ambiguity, reduce signalling overhead, and ensure consistent system behavior in deployment.
IV-F Integration with Emerging Use Cases
The adoption of any new waveform in 6G depends on its ability to reliably and efficiently support emerging use cases and service requirements. AFDM demonstrates strong potential in this regard, owing to its inherent flexibility, delay-Doppler robustness, and processing efficiency, well-aligned with key 6G paradigms.
For instance, AFDM is particularly well-suited for high-mobility scenarios such as V2X communications and SAGIN architectures involving low Earth orbit (LEO) constellations and unmanned aerial vehicles (UAVs), due to its resilience to Doppler shifts and delay spreads, supporting the vision of key verticals like automotive and future vehicular services [16].
In addition, the flexible design of AFDM supports advanced multifunctional applications, as shown by the previous discussions on the excellent integration of AFDM with ISAC framework, but also a plethora of other next-generation technologies such as semantic communications and over-the-air computing, both emerging enablers of future network intelligence.
Together, these capabilities position AFDM not only as an efficient transmission waveform, but as a flexible platform for addressing the diverse functional requirements of 6G.
IV-G Robustness to Practical Impairments
Robustness to real-world impairments remains a key requirement for waveform standardization, particularly for deployment in diverse hardware environments and under non-ideal channel conditions. Such impairments include carrier frequency offset (CFO), phase noise, amplifier nonlinearities, and quantization effects, all of which are known to degrade multicarrier waveforms like OFDM.
As evaluated, the AFDM exhibits natural resilience to these challenges, where the spreading in the DAFT domain distributes energy across time and frequency, mitigating sensitivity to CFO and providing improved phase noise tolerance. These properties position AFDM as a robust and scalable candidate for practical 6G scenarios where performance must be preserved despite front-end imperfections.
IV-H Testbeds, Trials, and Reference Designs
Finally, the transition from academic exploration to standardization necessitates rigorous validation through hardware testbeds, field trials, and open reference implementations. While AFDM has demonstrated strong theoretical performance across key metrics such as diversity gain, delay-Doppler resilience, and reduced ambiguity, its real-world viability remains to be fully established.
To this end, prototyping efforts based on software-defined radios and open-source platforms are essential to characterize implementation complexity, system-level performance, and compatibility with existing communication stacks. Such platforms can enable early experimentation with channel estimation, synchronization, and multi-user access techniques tailored to AFDM. These practical efforts will be instrumental in demonstrating the deployability of AFDM at scale and ensuring that its performance gains translate effectively into real-world environments, an essential step toward 6G adoption.
V Conclusion and Outlook
AFDM has emerged as a compelling candidate for the physical layer machinery of 6G systems, uniting robustness to delay-Doppler dispersion, full diversity, integrated sensing compatibility, and waveform-level flexibility within a single framework. The parametrized structure enables not only high communication performance but also opens new design frontiers for multi-functional operation.
Meanwhile, its backward compatibility, along with its readiness for frequency-domain processing, makes it particularly attractive for adoption in 6G, allowing cost-effective migration and coexistence in multi-RAT environments. Yet, the path to standardization remains open and requires focused development and discussions including bench-marking and link-and-system level demonstrations to further confirm our analysis.
All in all, with growing interest from both academia and industry, AFDM stands poised to help define the waveform landscape of the next generation of wireless standardization.